# Asymptotically Stable Quaternion-valued Hopfield-structured Neural Networks with Periodic Projection-based Supervised Learning Rules

### v1

@@ 1 Introduction
Hopfield neural network (HNN) could be recognized as the precursor of many fundamental models hnnsupport1 ; hnnsupport2 ; hnnsupport3 . It has a symmetric network topology and recurrent connections, becoming an attractor model with a dynamical structure, where neurons are activated and interact with each other at a certain frequency and tend to discrete equilibria. These equilibria and the attracting landscape in phase space embody the generalization capability of the neural network. We adopt the continuous time HNN HNN governed by the evolution Eq. (1.1) to expand and rebuild QSHNN in this research.
@@
In a more condensed vector form $\dot{\boldsymbol{v}}(t)=-\gamma\boldsymbol{v}(t)+\mu W\boldsymbol{\varphi}(\boldsymbol{v})+\mu\boldsymbol{b}$, Vector $\boldsymbol{v}$ represents the list of the neuron activations $\boldsymbol{v}_{j}$, matrix $W$ arranges the connection weights $\boldsymbol{w}_{ji}$ in matrix-vector multiplication, vector function $\boldsymbol{\varphi}$ imposes the activating function, typically hyperbolic tangent $\varphi(x)=(e^{x}-e^{-x})/(e^{x}+e^{-x})$, to each entry of vector $\boldsymbol{v}$. $\boldsymbol{b}$ serves as the bias vector and is significant to the representation ability of the network. Meaning of constant $\gamma,\mu$ in corresponds to the parameters of electronic components capacitance and resistor for the circuit implementation of classic HNN. The stability of HNNs is fundamental to their function as associative memory systems, enabling the convergence of network states to attractor patterns HNN . Earlier studies have extended HNNs to quaternion-valued systems, mainly under discrete-time formulations with split activation functions and fixed-point dynamics introduction ; qhnn1 . More recently, continuous-time modern HNNs have also been introduced allyouneed ; modernHNN , where the state evolution follows ordinary differential equations. However, these models remain largely confined to unsupervised paradigms driven by internal energy minimization, lacking explicit target tracking, structural control, or generalized learning principles.
@@
In contrast, our proposed model departs from this associative memory paradigm of HNN. We introduce a quaternion-valued supervised Hopfield-structured neural network (QSHNN), formulated as a continuous-time autonomous dynamical system. The system’s state evolves according to a quaternionic differential equation, and the network is trained to converge asymptotically to externally specified targets. The learning rule is derived analytically via Generalized $\mathbb{HR}$ (GHR) calculus GHR and incorporates a projection mechanism to preserve the block-wise quaternion structure of the weight matrix. GHR calculus has been established and verified as a complete and mature technique to expand the neural network over the quaternionic domain QB ; GHRsupport1 ; GHRsupport2 ; GHRsupport3 ; GHRsupport4 ; GhrSUPPORT5 , where the normal differential formulas are no longer valid because of the non-commutativity of quaternion algebra. This design guarantees both convergence and structural consistency.
@@
Although early Hopfield-type networks, including their quaternion-valued extensions, often rely on directly encoding the weight matrix using Hebbian or outer-product formulations, this approach suffers from several critical limitations. First, such weight constructions inherently lack scalability; they can only stably store a small number of target states before spurious attractors emerge or convergence fails lackA1 ; lackA2 ; lackC1 . Second, these methods assume all target patterns are fixed and known a priori, making the network unsuitable for tasks requiring adaptability, generalization, or dynamic reconfiguration lackB1 ; lackC1 . Third, direct encoding lacks an error-driven optimization mechanism, preventing the network from refining its behavior in response to task-specific objectives lackC1 ; lackC2 . As a result, while analytically convenient, direct weight specification severely restricts the practical applicability of Hopfield-type models in real-world control, learning, or representation settings. Unlike classical or previously proposed quaternion-valued Hopfield neural networks qvhnn1 ; qvhnn2 ; qhnn5 , our model operates under a supervised learning paradigm with continuous quaternion-valued trajectories, offering smooth, controllable dynamics suitable for robotics, trajectory generation, and feedback control systems. We position this work not as a generalization of existing QHNNs, but as the formulation of a new class of quaternionic neural dynamical systems with rigorous theoretical guarantees and task-driven behavior.
@@ 2 Background
The collection of all real $4\times 4$ matrices corresponding to quaternion left multiplication forms a smooth 4-dimensional embedded submanifold of $\mathbb{R}^{4\times 4}$, denoted by $\mathcal{L}$. This submanifold inherits a natural algebraic structure from $\mathbb{H}$ and constitutes a real matrix Lie group under composition matrixgroup . Specifically, $\mathcal{L}$ is isomorphic to the group $(\mathbb{H},\circ)$, and the group operation corresponds to matrix multiplication within $\mathcal{L}$. Each matrix $L(\boldsymbol{q})\in\mathcal{L}$ is uniquely determined by a quaternion $\boldsymbol{q}=s+x\boldsymbol{i}+y\boldsymbol{j}+z\boldsymbol{k}$, and the map $\boldsymbol{q}\mapsto L(\boldsymbol{q})$ is a group isomorphism. The tangent space at the identity element L(1) defines the associated Lie algebra, which consists of all real $4\times 4$ matrices that can be written as linear combinations of the infinitesimal generators $L(\boldsymbol{i}),L(\boldsymbol{j}),L(\boldsymbol{k})$.
@@
Definition 2.1 (Quaternion rotation rotation ).
@@
Definition 2.2 (Quaternion (left GHR) derivative GHR ).
@@
For the unknown solution to the evolution equation, which is of high-level nonlinearity and may not even exist in analytical form, we adopt Lyapunov theory for the analysis of stability. There are three different types of stability for a dynamical system that are in our consideration: Lyapunov stability, quasi-asymptotic stability, and asymptotic stability. The justification can be drawn from directly using the following Theorem (2.1).
Theorem 2.1 (Lyapunov Theorem Lyapunov ).
Consider the dynamical system of the form $\dot{\boldsymbol{x}}=\boldsymbol{f}(\boldsymbol{x},t)$ with the fixed point at the origin, where $\boldsymbol{f}(\boldsymbol{\boldsymbol{0},t})\equiv\boldsymbol{0}$. If there exists a positive definite scalar function $V(\boldsymbol{x},t)$, also named the Lyapunov energy function, such that the time derivative $\frac{dV}{dt}$ on the flow is semi-negative definite, as shown in Eq. (2.6), we have the zero solution being Lyapunov stable.
@@ 3 Methodology
The entire expression perfectly integrates quaternion embeddings instead of through complex mapping and splitting methods, where many required formalizations can be satisfied automatically. For higher order neuron systems with nonlinear activation functions, the solution would be too complicated and nonanalytic, shown in QDE1 ; QDE2 , and machine learning approaches get involved to make a qualitative control of this neural differential equation system.
@@
The learning of neural networks could follow the Hebbian learning rule hnnsupport3 like HNN, which is unsupervised, local, and physically inspired, different from constructing and minimizing the loss function, or the Delta learning rule delta1 ; delta2 which is supervised, global, and optimization-driven. By taking the differential of the sensitivity function, we derive the gradient direction in the weights parameter space for the loss function. This is the first stage of our model, for that the trained weights do not preserve quaternion structure, where we demand every $4\times 4$ block of the weights matrix is in the quaternion left multiplication submanifold. Hence we name the model supervised learning Hopfield-structured neural network (SHNN). The critical properties such as network stability and training accuracy are guaranteed with a rigorous mathematical mechanism through Thm. 3.1 and 3.2. For the second stage, we apply a technical skill called periodic projection. Every five iterations of strict gradient descent, we do a Frobenius orthogonal projection, as shown in Thm. (3.3), which may violate the principle of minimizing but imposes quaternionic structure on weight blocks. Globally, it will generate a periodic fluctuating path tending to the minimizer.
@@
where subscript $(\cdot)_{\varrho}$ is for quaternion components $\boldsymbol{q}=\boldsymbol{q}_{s}+\boldsymbol{q}_{x}\boldsymbol{i}+\boldsymbol{q}_{y}\boldsymbol{j}+\boldsymbol{q}_{z}\boldsymbol{k}$, and superscript is for quaternion involution introduced in Def. 2.1. This section shows that the quaternion structure-preserving descent exists, providing the foundation of the projection approach. But we would not apply this scheme for the algorithm. Formulas of quaternion calculus applied in this section could be found in GHR , especially the generalized chain rule deriving Eq. (3.15). These lead to the steepest descent direction, which is the gradient of the loss function $E(\boldsymbol{q^{*}})$:
@@ 5 Discussion
Based on the theoretical foundation of the proposed model, we plan to develop a complete application for robotic manipulator planning in the follow-up publication. Benchmark problems we consider are stated in Appendix E. Through these three levels of benchmarking, we can cover the widely recognized standard environment in the academic community and verify the versatility and scalability of direct sampling targets in industrial simulation, thereby fully demonstrating the combined advantages of QSHNN in terms of response speed, control accuracy, and online computing cost. Respectively, the baselines are the performance of current industrial algorithms, which are stated in Appendix F. In terms of the weaknesses mentioned of these baseline methods, QSHNN has the potential to outperform traditional strategies. Besides the guarantees on dynamical properties, powerful representation capability of QSHNN allows small-scale neural networks to complete tasks, thereby significantly improving the online planning reaction speed and reducing response time with the adjustment of the target.
@@
Throughout this research, we explore the potential to combine the memory-type Hopfield neural network with the mainstream method based on error propagation. The recent research on neuroscience revealed the surprising fact that the echo-location system of bats consists of only a small number of neurons as a core module bat , which motivates us to exploit the potential of bio-inspired recurrent neural networks. To implement this principle, we begin with the modification of continuous HNN and specify a concrete task in robotics by embedding physical information with quaternion. Though the theoretical foundation is established, only the development of the complete application with sufficient tests on the benchmark problems and comparison App. E with baselines App. F can make it more trustworthy, thereby giving us the confidence to further exploit the principle of general learning theory behind it, which is far more than the value of a single model.
@@ 6 Conclusions
We presented a quaternion-valued Hopfield-structured neural network that integrates structural consistency with smooth and stable learning dynamics. Modern HNN is mature in allyouneed , there are also many hypercomplex-valued versions qvhnn1 ; qvhnn2 ; qhnn5 , but the plasticity of supervision is usually weak, and the global stability is insufficient, where our model fills the gap in exploration in this aspect. By combining supervised training with periodic projection, the model preserves quaternionic structure while achieving accurate and reliable convergence. The bounded curvature of trajectories ensures smooth evolution, and the network’s design enables robust performance across randomly generated targets. Beyond orientation and control tasks, this approach also illustrates how embedding algebraic constraints into neural systems can yield both theoretical guarantees and practical benefits, offering a systematic path forward for structured learning in hypercomplex and non-commutative domains.
@@
Appendix A Proof of theorem 3.1: Existence and uniqueness of equilibrium
@@
Appendix B Proof of theorem 3.2: Lyapunov stability criterion
@@
Appendix C Proof of theorem 3.3: Projection to quaternion left multiplication manifold
@@ Appendix D Proof of smoothness: Curvature bound for the smoothness
Therefore, every component variable of $\boldsymbol{q}(t)$ has an upper bound of 4 on the second derivative. It gives a loose curvature bound on the trajectory in $\mathbb{R}^{n}$. In simpler terms, the path that q(t) traces through space is guaranteed to change direction smoothly. There are no sudden jerks, kinks, or high-frequency oscillations. The maximum bending or concavity of the trajectory is limited, which is important for smooth motion in control systems. If the actual engineering demands, we could get a more strict and smaller curvature $\kappa$ defined by curvature :
@@ Appendix E Benchmark problems
I. Robosuite PandaReach zhu2025robosuitemodularsimulationframework Robosuite’s Franka Panda Reach mission only requires end-to-target position and pose alignment, which is closer to the real industrial scene. The environment has 6 degrees of freedom + claws, allowing us to focus on evaluating attitude control rather than grabbing strategies.
 II. OpenAI Gym FetchReach plappert2018multigoalreinforcementlearningchallenging FetchReach is a 7-degree-of-freedom Fetch robotic arm target-alignment task. The target position is randomly generated in the environment (expandable to quaternions with attitude constraints), and the observation includes the relative position and direction of the end effector and the target, and the action directly gives the joint speed. In this environment, we can evaluate the response time, final error, and step time overhead of the QSHNN drive joint angle evolution to a specified quaternion pose.
 III. Random Workspace Target Alignment 6907751 Build a UR5 or KUKA LBR model directly in MuJoCo, uniformly sample the attitude targets (including position and quaternion direction) in their workspace, and test them on a large number of random initial-target pairs.

### v2

@@ 1 Introduction
Hopfield neural network (HNN) could be recognized as the precursor of many fundamental models Scellier (2021); Du (2011); Hopfield (1999). It has a symmetric network topology and recurrent connections, becoming an attractor model with a dynamical structure, where neurons are activated and interact with each other at a certain frequency and tend to discrete equilibria. These equilibria and the attracting landscape in phase space embody the generalization capability of the neural network. We adopt the continuous time HNN Hopfield (2007) governed by the evolution Eq. (1.1) to expand and rebuild QSHNN in this research.
@@
In a more condensed vector form $\dot{\boldsymbol{v}}(t)=-\gamma\boldsymbol{v}(t)+\mu W\boldsymbol{\varphi}(\boldsymbol{v})+\mu\boldsymbol{b}$, Vector $\boldsymbol{v}$ represents the list of the neuron activations $\boldsymbol{v}_{j}$, matrix $W$ arranges the connection weights $\boldsymbol{w}_{ji}$ in matrix-vector multiplication, vector function $\boldsymbol{\varphi}$ imposes the activating function, typically hyperbolic tangent $\varphi(x)=(e^{x}-e^{-x})/(e^{x}+e^{-x})$, to each entry of vector $\boldsymbol{v}$. $\boldsymbol{b}$ serves as the bias vector and is significant to the representation ability of the network. Meaning of constant $\gamma,\mu$ in corresponds to the parameters of electronic components capacitance and resistor for the circuit implementation of classic HNN. The stability of HNNs is fundamental to their function as associative memory systems, enabling the convergence of network states to attractor patterns Hopfield (2007). Earlier studies have extended HNNs to quaternion-valued systems, mainly under discrete-time formulations with split activation functions and fixed-point dynamics Greenblatt and Agaian (2018); Kobayashi,Masaki (2017). More recently, continuous-time modern HNNs have also been introduced Ramsauer et al. (2021); Achilli,Beatrice et al. (2025), where the state evolution follows ordinary differential equations. However, these models remain largely confined to unsupervised paradigms driven by internal energy minimization, lacking explicit target tracking, structural control, or generalized learning principles.
@@
In contrast, our proposed model departs from this associative memory paradigm of HNN. We introduce a quaternion-valued supervised Hopfield-structured neural network (QSHNN), formulated as a continuous-time autonomous dynamical system. The system’s state evolves according to a quaternionic differential equation, and the network is trained to converge asymptotically to externally specified targets. The learning rule is derived analytically via Generalized $\mathbb{HR}$ (GHR) calculus Xu et al. (2015) and incorporates a projection mechanism to preserve the block-wise quaternion structure of the weight matrix. GHR calculus has been established and verified as a complete and mature technique to expand the neural network over the quaternionic domain Pöppelbaum and Schwung (2022); Xu et al. (2017, 2016); Bourigault et al. (2024); Pöppelbaum and Schwung (2025); Takahashi et al. (2021), where the normal differential formulas are no longer valid because of the non-commutativity of quaternion algebra. This design guarantees both convergence and structural consistency.
@@
Although early Hopfield-type networks, including their quaternion-valued extensions, often rely on directly encoding the weight matrix using Hebbian or outer-product formulations, this approach suffers from several critical limitations. First, such weight constructions inherently lack scalability; they can only stably store a small number of target states before spurious attractors emerge or convergence fails Schmidgall et al. (2024); Athale and Stirk (1988); Ororbia et al. (2024). Second, these methods assume all target patterns are fixed and known a priori, making the network unsuitable for tasks requiring adaptability, generalization, or dynamic reconfiguration Alonso and Krichmar (2024); Ororbia et al. (2024). Third, direct encoding lacks an error-driven optimization mechanism, preventing the network from refining its behavior in response to task-specific objectives Ororbia et al. (2024); Launay et al. (2019). As a result, while analytically convenient, direct weight specification severely restricts the practical applicability of Hopfield-type models in real-world control, learning, or representation settings. Unlike classical or previously proposed quaternion-valued Hopfield neural networks Isokawa et al. (2008); Kobayashi,Masaki (2015); Valle (2014), our model operates under a supervised learning paradigm with continuous quaternion-valued trajectories, offering smooth, controllable dynamics suitable for robotics, trajectory generation, and feedback control systems. We position this work not as a generalization of existing QHNNs, but as the formulation of a new class of quaternionic neural dynamical systems with rigorous theoretical guarantees and task-driven behavior.
@@ 2 Background
The collection of all real $4\times 4$ matrices corresponding to quaternion left multiplication forms a smooth 4-dimensional embedded submanifold of $\mathbb{R}^{4\times 4}$, denoted by $\mathcal{L}$. This submanifold inherits a natural algebraic structure from $\mathbb{H}$ and constitutes a real matrix Lie group under composition Tapp (2010); Salamon (1986). Specifically, $\mathcal{L}$ is isomorphic to the group $(\mathbb{H},\circ)$, and the group operation corresponds to matrix multiplication within $\mathcal{L}$. Each matrix $L(\boldsymbol{q})\in\mathcal{L}$ is uniquely determined by a quaternion $\boldsymbol{q}=s+x\boldsymbol{i}+y\boldsymbol{j}+z\boldsymbol{k}$, and the map $\boldsymbol{q}\mapsto L(\boldsymbol{q})$ is a group isomorphism. The tangent space at the identity element L(1) defines the associated Lie algebra, which consists of all real $4\times 4$ matrices that can be written as linear combinations of the infinitesimal generators $L(\boldsymbol{i}),L(\boldsymbol{j}),L(\boldsymbol{k})$.
@@
Definition 2.1 (Quaternion rotation Ward (2012)).
@@
Definition 2.2 (Quaternion (left GHR) derivative Xu et al. (2015)).
@@
For the unknown solution to the evolution equation, which is of high-level nonlinearity and may not even exist in analytical form, we adopt Lyapunov theory for the analysis of stability. There are three different types of stability for a dynamical system that are in our consideration: Lyapunov stability, quasi-asymptotic stability, and asymptotic stability. The justification can be drawn from directly using the following Thm. (2.1).
Theorem 2.1 (Lyapunov Theorem Sastry (1999)).
Consider the dynamical system of the form $\dot{\boldsymbol{x}}=\boldsymbol{f}(\boldsymbol{x},t)$ with the fixed point at the origin, where $\boldsymbol{f}(\boldsymbol{0},t)\equiv\boldsymbol{0}$. If there exists a positive definite scalar function $V(\boldsymbol{x},t)$, also named the Lyapunov energy function, such that the time derivative $\frac{dV}{dt}$ on the flow is semi-negative definite, as shown in Eq. (2.6), we have the zero equilibrium being Lyapunov stable.
@@ 3 Methodology
The entire expression perfectly integrates quaternion embeddings instead of through complex mapping and splitting methods, where many required formalizations can be satisfied automatically. For higher order neuron systems with nonlinear activation functions, the solution would be too complicated and nonanalytic, shown in Lv et al. (2023); Kou and Xia (2018), and machine learning approaches get involved to make a qualitative control of this neural differential equation system.
@@
The learning of neural networks could follow the Hebbian learning rule Hopfield (1999) like HNN, which is unsupervised, local, and physically inspired, different from constructing and minimizing the loss function, or the Delta learning rule Rumelhart et al. (1986); Widrow and Lehr (1990) which is supervised, global, and optimization-driven. By taking the differential of the sensitivity function, we derive the gradient direction in the weights parameter space for the loss function. This is the first stage of our model, for that the trained weights do not preserve quaternion structure, where we demand every $4\times 4$ block of the weights matrix is in the quaternion left multiplication submanifold. Hence we name the model supervised learning Hopfield-structured neural network (SHNN). The critical properties such as network stability and training accuracy are guaranteed with a rigorous mathematical mechanism through Thm. 3.1 and 3.2. For the second stage, we apply a technical skill called periodic projection. Every five iterations of strict gradient descent, we do a Frobenius orthogonal projection, as shown in Thm. (3.3), which may violate the principle of minimizing but imposes quaternionic structure on weight blocks. Globally, it will generate a periodic fluctuating path tending to the minimizer.
@@
where subscript $(\cdot)_{\varrho}$ is for quaternion components $\boldsymbol{q}=\boldsymbol{q}_{s}+\boldsymbol{q}_{x}\boldsymbol{i}+\boldsymbol{q}_{y}\boldsymbol{j}+\boldsymbol{q}_{z}\boldsymbol{k}$, and superscript is for quaternion involution introduced in Def. 2.1. This section shows that the quaternion structure-preserving descent exists, providing the foundation of the projection approach. But we would not apply this scheme for the algorithm. Formulas of quaternion calculus applied in this section could be found in Xu et al. (2015), especially the generalized chain rule deriving Eq. (3.15). These lead to the steepest descent direction, which is the gradient of the loss function $E(\boldsymbol{q^{*}})$:
@@ 5 Discussion
Based on the theoretical foundation of the proposed model, we plan to develop a complete application for robotic manipulator planning in the follow-up publication. Benchmark problems we consider are stated in App. E. Through these three levels of benchmarking, we can cover the widely recognized standard environment in the academic community and verify the versatility and scalability of direct sampling targets in industrial simulation, thereby fully demonstrating the combined advantages of QSHNN in terms of response speed, control accuracy, and online computing cost. Respectively, the baselines are the performance of current industrial algorithms, which are stated in App. F. In terms of the weaknesses mentioned of these baseline methods, QSHNN has the potential to outperform traditional strategies. Besides the guarantees on dynamical properties, powerful representation capability of QSHNN allows small-scale neural networks to complete tasks, thereby significantly improving the online planning reaction speed and reducing response time with the adjustment of the target.
@@
Throughout this research, we explore the potential to combine the memory-type Hopfield neural network with the mainstream method based on error propagation. The recent research on neuroscience revealed the surprising fact that the echo-location system of bats consists of only a small number of neurons as a core module Forli et al. (2025), which motivates us to exploit the potential of bio-inspired recurrent neural networks. To implement this principle, we begin with the modification of continuous HNN and specify a concrete task in robotics by embedding physical information with quaternion. Though the theoretical foundation is established, only the development of the complete application with sufficient tests on the benchmark problems and comparison App. E with baselines App. F can make it more trustworthy, thereby giving us the confidence to further exploit the principle of general learning theory behind it, which is far more than the value of a single model.
@@ 6 Conclusions
We presented a quaternion-valued Hopfield-structured neural network that integrates structural consistency with smooth and stable learning dynamics. Modern HNN is mature in Ramsauer et al. (2021), there are also many hypercomplex-valued versions Isokawa et al. (2008); Kobayashi,Masaki (2015); Valle (2014), but the plasticity of supervision is usually weak, and the global stability is insufficient, where our model fills the gap in exploration in this aspect. By combining supervised training with periodic projection, the model preserves quaternionic structure while achieving accurate and reliable convergence. The bounded curvature of trajectories ensures smooth evolution, and the network’s design enables robust performance across randomly generated targets. Beyond orientation and control tasks, this approach also illustrates how embedding algebraic constraints into neural systems can yield both theoretical guarantees and practical benefits, offering a systematic path forward for structured learning in hypercomplex and non-commutative domains.
@@
Appendix A Proof of Thm. 3.1: Existence and uniqueness of equilibrium
@@
Appendix B Proof of Thm. 3.2: Lyapunov stability criterion
@@
Appendix C Proof of Thm. 3.3: Projection to quaternion left multiplication manifold
@@ Appendix D Proof of smoothness: Curvature bound for the smoothness
Therefore, every component variable of $\boldsymbol{q}(t)$ has an upper bound of 4 on the second derivative. It gives a loose curvature bound on the trajectory in $\mathbb{R}^{n}$. In simpler terms, the path that q(t) traces through space is guaranteed to change direction smoothly. There are no sudden jerks, kinks, or high-frequency oscillations. The maximum bending or concavity of the trajectory is limited, which is important for smooth motion in control systems. If the actual engineering demands, we could get a more strict and smaller curvature $\kappa$ defined by [5]:
@@ Appendix E Benchmark problems
I. Robosuite PandaReach [39] Robosuite’s Franka Panda Reach mission only requires end-to-target position and pose alignment, which is closer to the real industrial scene. The environment has 6 degrees of freedom + claws, allowing us to focus on evaluating attitude control rather than grabbing strategies.
 II. OpenAI Gym FetchReach [21] FetchReach is a 7-degree-of-freedom Fetch robotic arm target-alignment task. The target position is randomly generated in the environment (expandable to quaternions with attitude constraints), and the observation includes the relative position and direction of the end effector and the target, and the action directly gives the joint speed. In this environment, we can evaluate the response time, final error, and step time overhead of the QSHNN drive joint angle evolution to a specified quaternion pose.
 III. Random Workspace Target Alignment [32] Build a UR5 or KUKA LBR model directly in MuJoCo, uniformly sample the attitude targets (including position and quaternion direction) in their workspace, and test them on a large number of random initial-target pairs.